OpenStudy (anonymous):

hi everyone Solve -4(x + 10) - 6 = -3(x - 2). A) -40 B) -46 C) -52 D) 52

OpenStudy (ageta):

Let's solve your equation step-by-step. −4(x+10)−6=−3(x−2) Step 1: Simplify both sides of the equation. −4x−46=−3x+6 Step 2: Add 3x to both sides. −4x−46+3x=−3x+6+3x −x−46=6 Step 3: Add 46 to both sides. −x−46+46=6+46 −x=52 Step 4: Divide both sides by -1. −x −1 = 52 −1 x=−52 Answer: x=−52

OpenStudy (uri):

-4(x + 10) - 6 = -3(x - 2). -4*x +4*10 -6 =-3*x+3*2 Let's do step by step,shall we? @extremearianna

OpenStudy (ageta):

the answer is -52

OpenStudy (anonymous):

Thanks guys, I realize this is a lot easier than I thought.. I just get my order of operations confused sometimes..

OpenStudy (uri):

...okay.

OpenStudy (ageta):

no problem

OpenStudy (12man):

your welcome